# Self-consistency error correction for accurate machine learning potentials from variational Monte Carlo

### v1

@@ 1 Introduction
Since their introduction 1, 2, machine learning interatomic potentials (MLIPs) have revolutionized the field of molecular dynamics (MD) 3, 4, 5, 6, 7, 8, 9, 10, 11, 12. These models can learn the potential energy surface (PES) of a given system starting from a representative set of configurations, called the training set. Each element of this set is labeled by its physical properties (e.g., energy, ionic forces, pressure, etc.), usually computed using ab initio electronic structure methods. After it is trained, an MLIP is capable of reproducing the observables of the system on new configurations with the accuracy of the original method but with a computational cost of several orders of magnitude smaller 2. This has opened the possibility of running MD simulations at ab initio accuracy with sizes and lengths that were previously out of reach 13. By far, the most common method used for generating MLIPs datasets is density functional theory (DFT) 14, which can easily produce $\sim 10^{4}$-$10^{5}$ training points with modest computational resources, usually sufficient to train data-heavy machine learning frameworks 7, 15. For several systems, however, DFT fails to capture the strong effect of electronic correlations or may provide results that strongly depend on the choice of the exchange-correlation functional. Quantum Monte Carlo (QMC) methods are a series of stochastic algorithms that provide state-of-the-art results for a large class of systems, from molecules to solids 16, 17. Although more accurate, QMC approaches have a much larger computational cost than DFT. In the context of MLIPs, this implies that the number of training configurations that can be generated within a given amount of resources is generally smaller than the one attainable with mean field methods. Recently proposed solutions to this problem make use of data-efficient schemes in combination with hierarchical machine learning (or $\Delta$-learning 18) to construct QMC-based MLIPs 19, 20. However, within the most popular QMC approach, i.e., variational Monte Carlo (VMC), another issue prevents a straightforward application of MLIPs, presented in what follows.
@@
where $\exp\left(J\right)$ is a bosonic term (i.e., symmetric under particle exchange) called the Jastrow factor, and $\Phi_{\textrm{AS}}$ is an antisymmetric (AS) part. Both terms contain variational parameters, whose number can lead to flexible WFs. However, there is a practical difficulty in optimizing a large number of variables in a stochastic setting, particularly those belonging to the AS part. Thus, only the Jastrow parameters are usually optimized, and $\Phi_{\textrm{AS}}\equiv\Phi_{\textrm{SD}}$ is often taken as a single Slater determinant (SD) obtained from a DFT calculation 17. This gives rise to the frozen-orbital Jastrow Slater determinant (JSD) wave function, an appealing candidate for VMC-based machine learning applications, since it provides an excellent tradeoff between accuracy and computational cost. Indeed, the Jastrow optimization is much less demanding than the full optimization also involving the determinant. Nevertheless, the use of the frozen Slater determinant introduces a bias in the VMC forces and pressures computed with this WF. This bias has been dubbed as the self-consistency error  21, 22, and we are going to show that it is particularly relevant in the context of VMC-based MLIPs. Even though machine learning potentials are, by definition, consistent, a biased training set can nonetheless spoil the final accuracy of the model. A solution would be to use only the energies for the training; this, however, usually increases by orders of magnitude the number of configurations necessary to reach a given accuracy 23, 24.
@@
In Ref. 22, a suitable correction was developed to remove the self-consistency error, adding a relatively small computational overhead to the calculation. The aim of this work is to show the actual impact of the self-consistency error on an MLIP in a working case scenario, and to demonstrate how its correction can be applied to generate unbiased VMC datasets with a JSD WF for machine-learning applications. As test case, we consider high-pressure hydrogen, a widely studied system with applications ranging from planet modeling 25 to inertial confinement fusion 26, 27, 28, and often used to benchmark numerical methods 29. In particular, we focus on density and temperatures relevant for the Hugoniot curve, i.e., the set of possible states reachable with a shock wave 30. The deuterium Hugoniot was recently computed in Ref. 20 using MLIPs trained on accurate QMC data. There, the self-consistency error was avoided by relaxing the full set of variational parameters in the WF, i.e., including both the Jastrow factor and the antisymmetric part. The resulting optimization improves the consistency of the forces and pressures and also changes the underlying PES, at a computational cost about 10 times larger than the standard frozen-orbital JSD optimization.
@@
Here, we explore a different path, by showing that a very similar accuracy can be obtained by training models on a frozen-orbital JSD dataset supplemented by the self-consistency error correction. In particular, we compute VMC forces and pressures on the same configurations using a JSD WF, with and without corrections, and use them to train the different models. The performances of these MLIPs are compared with the one of an MLIP trained on the fully optimized dataset of Ref. 20, which we take as the reference. In this way, we show the relevance of the self-consistency error in the MLIP generation and the need for its correction, by studying the impact of both the correction and the full optimization of the WF on thermodynamic equilibrium quantities such as pressure and radial distribution functions. This work demonstrates the importance of obtaining unbiased QMC datasets for MLIPs at the VMC level using cheap-to-optimize JSD-type WFs, including also ionic forces and stress components.
@@ 2 Methods
We briefly present the main methods used in this work. In Sec. 2.1 we introduce general aspects of MLIPs, focusing on one specific framework, i.e., the kernel ridge regression, and on the loss function used for model training. In Sec.2.2, we review the forces and pressure correction developed in Ref. 22 to solve the self-consistency error in VMC and describe its implementation in our WF representation.
@@ 2.1 Generalities of MLIPs
The basic goal of any MLIPs is to predict the energy of a given $N$-atom configuration described by its ionic positions, i.e., $E=E\left(\mathbf{R}_{1},\cdots,\mathbf{R}_{N}\right)$. We now suppose that $E$ can be expressed as a sum of atomic contributions 2, each depending on the relative coordinates of all the other atoms with respect to the $i$-th one:
@@
Here, we refer to $\mathcal{R}_{i}$ as the local environment around the $i$-th atom and to $e\left(\mathcal{R}_{i}\right)$ as the local atomic energy, and suppose the minimum image convention holds when periodic boundary conditions are applied. Often, the function $e\left(\mathcal{R}_{i}\right)$ in Eq.(1) is considered local, i.e., only dependent on the atoms closer than a certain cutoff radius $r_{c}$. Different types of MLIPs may be distinguished by their internal representation of the local atomic energy. Many techniques have been successfully applied over the years, including deep neural networks 2, 4, 11, Gaussian approximation potentials 31, 32, kernel ridge regression 33, 19, graph neural networks 34, 35, and more. Independently of the specific architecture, $e\left(\mathcal{R}_{i}\right)$ usually depends on a set of model parameters $\mathbf{p}$ which have to be optimized:
@@
where $\mathcal{R}_{\nu}$, $\nu=1,\dots,N_{\textrm{env}}$ belong to a subset of all the local environments in the training set, and $K(\mathcal{R},\mathcal{R_{\nu}})$ is the normalized kernel between $\mathcal{R}$ and $\mathcal{R}_{\nu}$. Following Refs 19 and 20, we adopt a kernel based on a variant of the Smooth Overlap of Atomic Positions (SOAP) descriptor 36, 37(see App. A.1), and we select the $N_{\textrm{env}}$ local environments using the farthest point sampling method 19, according to the distance introduced by the kernel $K$. Notice how plugging Eq.(4) into the loss function (3) and taking the derivatives with respect to the variational coefficients turns the minimization problem into a linear system, which is then solved by the conjugate gradient method. We conclude this Section by stressing the fact that, by definition, any MLIPs will give energy derivatives (e.g., forces and virial stress components) that are consistent with their underlying potential energy surface. Any effect of potential inconsistency in the dataset, such as the ones studied in this work, will manifest in the final performance of the models and its dependence on the relative training weights $W_{E}$, $W_{F}$, and $W_{P}$. This is demonstrated in Sec.3.2, where we compare the quality of MLIPs trained on different datasets as a function of $W_{E}$, $W_{F}$, and $W_{P}$.
@@ 2.2 Self-consistency error in VMC and its correction
[…]
@@
Discarding the term in Eq.(6) causes an inconsistency between the forces and the underlying PES computed with a JSD WF, meaning that $\mathbf{F}^{\textrm{VMC}}_{i}\neq-\frac{\mathop{dE}}{\mathop{d\mathbf{R}_{i}}}$. This is called the self-consistency error (SCE) 21, 22. An analogous discussion applies to other energy derivatives, such as the virial pressure, that contain a similar bias and for which we can derive similar corrections.
In our applications, following the implementation of the TurboRVB package 45 (here used for all the QMC calculations), we consider a representation of the SD part in terms of the antisymmetrized geminal power (AGP)
@@
In our case, the AOs are taken as Gaussian type orbitals (GTOs). For periodic systems, such as the ones considered in this work, the localized basis is appropriately generalized to satisfy periodic or twisted boundary conditions 46, depending on the phase $\theta$ ($\theta^{\prime}$) for spin-up (spin-down) electrons. Here, $\theta=-\theta^{\prime}$, such that the global phase factor is canceled, making the function $f$ invariant under a global translation. Fixing this gauge leads to a numerically stable evaluation of the $d\lambda_{k,k^{\prime}}/d\mathbf{R}$ derivatives implied by Eq. 6.
An alternative expression for the function $f$ can also be obtained using molecular orbitals (MOs) $\Phi_{n}(\mathbf{r})=\sum_{k}c_{n,k}\phi_{k}(\mathbf{r})$. In particular, when the number $M$ of MOs is equal to $N/2$, the resulting AGP is exactly equivalent to a single SD, i.e., $\Phi_{\textrm{AGP}}\equiv\Phi_{\textrm{SD}}$. Within TurboRVB, the MOs (and the corresponding $\lambda_{k,k^{\prime}}$ matrix) are initialized by a local density approximation 47 (LDA) DFT calculation. This is performed in a basis consisting of the aforementioned AOs with an additional one-body Jastrow factor to automatically take into account Kato cusp conditions 48, 45. Because of this, the electronic integrals entering into the Kohn-Sham equations are evaluated on a real space grid with a suitably chosen lattice space. Finally, we also mention that the basis is regularized by removing any eigenvector of the overlap matrix $S_{k,k^{\prime}}=\langle\phi_{k}|\phi_{k^{\prime}}\rangle$ with a corresponding eigenvalue ($s_{i}$) satisfying the condition $s_{i}/s_{\rm max}<\epsilon_{\textrm{cut}}$ 49, where $s_{\rm max}$ is the maximum eigenvalue and $\epsilon_{\textrm{cut}}$ is set to $10^{-7}$ in this study.
@@
where $\mathcal{O}_{k,k^{\prime}}=\frac{\partial\log\Psi}{\partial\lambda_{k,k^{\prime}}}$ and $\overline{\mathcal{O}}_{k,k^{\prime}}=\langle\mathcal{O}_{k,k^{\prime}}\rangle$. In the previous expression, the logarithmic derivatives in $\mathcal{O}_{k,k^{\prime}}$ are efficiently computed using the adjoint algorithmic differentiation 42, while the derivatives of the parameters, $\frac{d\lambda_{k,k^{\prime}}}{d\mathbf{R}_{i}}$, are evaluated within DFT-LDA using the finite differences method (FDM), as described in Ref. 22. For the virial isotropic pressure, an analogous derivation yields the following expression for the correction term:
@@ 3 Results
Henceforth, we analyze how the SCE affects the training of MLIPs and discuss how the correction described in Sec. 2.2 can effectively be implemented to improve the accuracy of the resulting models. With this aim, we consider a dataset made of pristine hydrogen configurations for which we computed VMC energy, forces, and pressure using different types of WF and with/without bias correction. The dataset is the same as the one used in Ref. 20 and comprises $558$ configurations of $128$ atoms each. These configurations were extracted from DFT-driven MD simulations at temperatures between $4000$ K and $20000$ K and Wigner-Seitz radii $r_{s}$ between $1.80$ and $2.12$. This range of thermodynamic conditions corresponds to the expected location of the Hugoniot curve of deuterium, defined as the set of states that can be reached by producing a shock wave in a sample. For these configurations, we fully optimized the wave function, going beyond the frozen-orbital JSD ansatz. As previously mentioned, we take this dataset as the reference one. In Sec. 3.1, we study the SCE for the same set of configurations as the reference dataset, and show the effect of both forces and pressure corrections. In Sec. 3.2 we directly compare the accuracy of MLIPs trained on a biased dataset (i.e. affected by the SCE) and a corrected one, respectively. We also compare both models with the reference dataset used in Ref. 20. Finally, in Sec. 3.3, we analyze the results of molecular dynamics simulations driven by the different MLIPs.

### v2

@@ 1 Introduction
Since their introduction Blank1995, Behler2007, machine learning interatomic potentials (MLIPs) have revolutionized the field of molecular dynamics (MD) Handley2010, Behler2011, Manzhos2015, Botu2017, Deringer2019, Wood2019, Dral2020, Noé2020, Behler2021, Unke2021. These models can learn the potential energy surface (PES) of a given system starting from a representative set of configurations, called the training set. Each element of this set is labeled by its physical properties (e.g., energy, ionic forces, pressure, etc.), usually computed using ab initio electronic structure methods. After it is trained, an MLIP is capable of reproducing the observables of the system on new configurations with the accuracy of the original method but with a computational cost of several orders of magnitude smaller Behler2007. This has opened the possibility of running MD simulations at ab initio accuracy with sizes and lengths that were previously out of reach Jia2020. By far, the most common method used for generating MLIPs datasets is density functional theory (DFT) Martin2004, which can easily produce $\sim 10^{4}$-$10^{5}$ training points with modest computational resources, usually sufficient to train data-heavy machine learning frameworks Deringer2019, Schleder2019. For several systems, however, DFT fails to capture the strong effect of electronic correlations or may provide results that strongly depend on the choice of the exchange-correlation functional. Quantum chemistry methods, such as the density matrix renormalization group (DMRG) Olivares-Amaya2015, and the complete active space self-consistent field (CASSCF)Jensen1987, Werner1985, provide a much better description of correlation, but their applicability to periodic extended systems is limited by their large computational cost. Because of this, even if significant progress has been made recently, also thanks to the implementation of new machine learning schemes Rath2025, the range of applicability of these approaches remains limited to a few tens of atoms and a few $ps$-long trajectories. Quantum Monte Carlo (QMC) methods are a series of stochastic algorithms that provide state-of-the-art results for a large class of systems, from molecules to solids Foulkes2001, Becca2017, while being less computationally demanding. Despite this, QMC approaches still have a cost about 10-100 times larger than DFT. In the context of MLIPs, this implies that the number of training configurations that can be generated within a given amount of resources is generally smaller than the one attainable with mean-field methods. Recently proposed solutions to this problem make use of data-efficient schemes in combination with hierarchical machine learning (or $\Delta$-learning Ramakrishnan2015) to construct QMC-based MLIPs Tirelli2022, Tenti2024. However, within the most popular QMC approach, i.e., variational Monte Carlo (VMC), another issue prevents a straightforward application of MLIPs, presented in what follows.
@@
where $\exp\left(J\right)$ is a bosonic term (i.e., symmetric under particle exchange) called the Jastrow factor, and $\Phi_{\textrm{AS}}$ is an antisymmetric (AS) part. Both terms contain variational parameters, whose number can lead to flexible WFs. However, there is a practical difficulty in optimizing a large number of variables in a stochastic setting, particularly those belonging to the AS part. Thus, only the Jastrow parameters are usually optimized, and $\Phi_{\textrm{AS}}\equiv\Phi_{\textrm{SD}}$ is often taken as a single Slater determinant (SD) obtained from a DFT calculation Becca2017. This gives rise to the frozen-orbital Jastrow Slater determinant (JSD) wave function, an appealing candidate for VMC-based machine learning applications, since it provides an excellent tradeoff between accuracy and computational cost. Indeed, the Jastrow optimization is much less demanding than the full optimization also involving the determinant. Nevertheless, the use of the frozen Slater determinant introduces a bias in the VMC forces and pressures computed with this WF. This bias has been dubbed as the self-consistency error  Tiihonen2021, Nakano2024, and we are going to show that it is particularly relevant in the context of VMC-based MLIPs. Even though machine learning potentials are, by definition, consistent, a biased training set can nonetheless spoil the final accuracy of the model. A solution would be to use only the energies for the training; this, however, usually increases by orders of magnitude the number of configurations necessary to reach a given accuracy Chmiela2017, Christensen2020.
@@
In Ref. Nakano2024, a suitable correction was developed to remove the self-consistency error, adding a relatively small computational overhead to the calculation. The aim of this work is to show the actual impact of the self-consistency error on an MLIP in a working case scenario, and to demonstrate how its correction can be applied to generate unbiased VMC datasets with a JSD WF for machine-learning applications. As a test case, we consider high-pressure hydrogen, a widely studied system with applications ranging from planet modeling Helled2020 to inertial confinement fusion Hu2011, Craxton2015, Hu2015, and often used to benchmark numerical methods Bonitz2024. It is also a system where DFT usually fails, when compared with the experimental dataMonacelli2023, making it an ideal testbed for our purposes. In particular, we focus on density and temperatures relevant for the Hugoniot curve, i.e., the set of possible states reachable with a shock wave Duvall1977. The deuterium Hugoniot was recently computed in Ref. Tenti2024 using MLIPs trained on accurate QMC data. There, the self-consistency error was avoided by relaxing the full set of variational parameters in the WF, i.e., including both the Jastrow factor and the antisymmetric part. The resulting optimization improves the consistency of the forces and pressures and also changes the underlying PES, at a computational cost about 10 times larger than the standard frozen-orbital JSD optimization.
@@
Here, we explore a different path, by showing that a very similar accuracy can be obtained by training models on a frozen-orbital JSD dataset supplemented by the self-consistency error correction. In particular, we compute VMC forces and pressures on the same configurations using a JSD WF, with and without corrections, and use them to train the different models. The performances of these MLIPs are compared with the one of an MLIP trained on the fully optimized dataset of Ref. Tenti2024, which we take as the reference. In this way, we show the relevance of the self-consistency error in the MLIP generation and the need for its correction, by studying the impact of both the correction and the full optimization of the WF on thermodynamic equilibrium quantities such as pressure and radial distribution functions. This work demonstrates the importance of obtaining unbiased QMC datasets for MLIPs at the VMC level using cheap-to-optimize JSD-type WFs, including also ionic forces and stress components.
@@ 2 Methods
We briefly present the main methods used in this work. In Sec. 2.1 we introduce general aspects of MLIPs, focusing on one specific framework, i.e., the kernel ridge regression, and on the loss function used for model training. In Sec.2.2, we review the forces and pressure correction developed in Ref. Nakano2024 to solve the self-consistency error in VMC and describe its implementation in our WF representation.
@@ 2.1 Generalities of MLIPs
The basic goal of any MLIPs is to predict the energy of a given $N$-atom configuration described by its ionic positions, i.e., $E=E\left(\mathbf{R}_{1},\cdots,\mathbf{R}_{N}\right)$. We now suppose that $E$ can be expressed as a sum of atomic contributions Behler2007, each depending on the relative coordinates of all the other atoms with respect to the $i$-th one:
@@
Here, we refer to $\mathcal{R}_{i}$ as the local environment around the $i$-th atom and to $e\left(\mathcal{R}_{i}\right)$ as the local atomic energy, and suppose the minimum image convention holds when periodic boundary conditions are applied. Often, the function $e\left(\mathcal{R}_{i}\right)$ in Eq.(1) is considered local, i.e., only dependent on the atoms closer than a certain cutoff radius $r_{c}$. Different types of MLIPs may be distinguished by their internal representation of the local atomic energy. Many techniques have been successfully applied over the years, including deep neural networks Behler2007, Behler2011, Behler2021, Gaussian approximation potentials Bartok2010, Bartok2015, kernel ridge regression Rupp2012, Tirelli2022, graph neural networks Batzner2022, Batatia2022, and more. Independently of the specific architecture, $e\left(\mathcal{R}_{i}\right)$ usually depends on a set of model parameters $\mathbf{p}$ which have to be optimized:
@@
where $\mathcal{R}_{\nu}$, $\nu=1,\dots,N_{\textrm{env}}$ belong to a subset of all the local environments in the training set, and $K(\mathcal{R},\mathcal{R_{\nu}})$ is the normalized kernel between $\mathcal{R}$ and $\mathcal{R}_{\nu}$. Following Refs Tirelli2022 and Tenti2024, we adopt a kernel based on a variant of the Smooth Overlap of Atomic Positions (SOAP) descriptor Bartok2013, De2016(see App. A.1), and we select the $N_{\textrm{env}}$ local environments using the farthest point sampling method Tirelli2022, according to the distance introduced by the kernel $K$. Notice how plugging Eq.(4) into the loss function (3) and taking the derivatives with respect to the variational coefficients turns the minimization problem into a linear system, which is then solved by the conjugate gradient method. We conclude this Section by stressing the fact that, by definition, any MLIPs will give energy derivatives (e.g., forces and virial stress components) that are consistent with their underlying potential energy surface. Any effect of potential inconsistency in the dataset, such as the ones studied in this work, will manifest in the final performance of the models and its dependence on the relative training weights $W_{E}$, $W_{F}$, and $W_{P}$. This is demonstrated in Sec.3.2, where we compare the quality of MLIPs trained on different datasets as a function of $W_{E}$, $W_{F}$, and $W_{P}$.
@@ 2.2 Self-consistency error in VMC and its correction
where $\langle\cdots\rangle$ indicates the quantum expectation value on the wave function $\Psi$ evaluated stochastically, $E_{L}\equiv\hat{H}\Psi/\Psi$ the local energy with $\hat{H}$ being the Hamiltonian of the system and $E=\langle E_{L}\rangle$. To obtain an efficient and statistically meaningful value of the force in Eq. (5) with finite variance, techniques such as the zero-variance zero-bias principle Assaraf2003, the space warp transformation Umrigar1989 and reweighting Assaraf2000, Assaraf2003, Attaccalite2008, Sorella2010, Filippi2016, Nakano2022 are applied. $\mathbf{F}^{\textrm{VMC}}_{i}$ is unbiased only when all the variational parameters of the WF are at their variational minimum. As mentioned earlier, this assumption is not true for partially optimized WFs, such as the frozen-orbital JSD, one of the most popular choices. In this case, an extra term
@@
Discarding the term in Eq.(6) causes an inconsistency between the forces and the underlying PES computed with a JSD WF, meaning that $\mathbf{F}^{\textrm{VMC}}_{i}\neq-\frac{\mathop{dE}}{\mathop{d\mathbf{R}_{i}}}$. This is called the self-consistency error (SCE) Tiihonen2021, Nakano2024. An analogous discussion applies to other energy derivatives, such as the virial pressure, that contain a similar bias and for which we can derive similar corrections.
In our applications, following the implementation of the TurboRVB package Nakano2020 (here used for all the QMC calculations), we consider a representation of the SD part in terms of the antisymmetrized geminal power (AGP)
@@
In our case, the AOs are taken as Gaussian type orbitals (GTOs). For periodic systems, such as the ones considered in this work, the localized basis is appropriately generalized to satisfy periodic or twisted boundary conditions Dovesi2018, depending on the phase $\theta$ ($\theta^{\prime}$) for spin-up (spin-down) electrons. Here, $\theta=-\theta^{\prime}$, such that the global phase factor is canceled, making the function $f$ invariant under a global translation. Fixing this gauge leads to a numerically stable evaluation of the $d\lambda_{k,k^{\prime}}/d\mathbf{R}$ derivatives implied by Eq. 6.
An alternative expression for the function $f$ can also be obtained using molecular orbitals (MOs) $\Phi_{n}(\mathbf{r})=\sum_{k}c_{n,k}\phi_{k}(\mathbf{r})$. In particular, when the number $M$ of MOs is equal to $N/2$, the resulting AGP is exactly equivalent to a single SD, i.e., $\Phi_{\textrm{AGP}}\equiv\Phi_{\textrm{SD}}$. Within TurboRVB, the MOs (and the corresponding $\lambda_{k,k^{\prime}}$ matrix) are initialized by a local density approximation Perdew1981 (LDA) DFT calculation. This is performed in a basis consisting of the aforementioned AOs with an additional one-body Jastrow factor to automatically take into account Kato cusp conditions Kato1957, Nakano2020. Because of this, the electronic integrals entering into the Kohn-Sham equations are evaluated on a real space grid with a suitably chosen lattice space. Finally, we also mention that the basis is regularized by removing any eigenvector of the overlap matrix $S_{k,k^{\prime}}=\langle\phi_{k}|\phi_{k^{\prime}}\rangle$ with a corresponding eigenvalue ($s_{i}$) satisfying the condition $s_{i}/s_{\rm max}<\epsilon_{\textrm{cut}}$ Azadi2010, where $s_{\rm max}$ is the maximum eigenvalue and $\epsilon_{\textrm{cut}}$ is set to $10^{-7}$ in this study. This regularization is very general and is also able to automatically cure the problem of basis-set redundancy coming from delocalized periodic Gaussians.
@@
where $\mathcal{O}_{k,k^{\prime}}=\frac{\partial\log\Psi}{\partial\lambda_{k,k^{\prime}}}$ and $\overline{\mathcal{O}}_{k,k^{\prime}}=\langle\mathcal{O}_{k,k^{\prime}}\rangle$. In the previous expression, the logarithmic derivatives in $\mathcal{O}_{k,k^{\prime}}$ are efficiently computed using the adjoint algorithmic differentiation Sorella2010, while the derivatives of the parameters, $\frac{d\lambda_{k,k^{\prime}}}{d\mathbf{R}_{i}}$, are evaluated within DFT-LDA using the finite differences method (FDM), as described in Ref. Nakano2024. For the virial isotropic pressure, an analogous derivation yields the following expression for the correction term:
@@
The above Equation can be easily extended to the non-isotropic case.
@@ 3 Results
Henceforth, we analyze how the SCE affects the training of MLIPs and discuss how the correction described in Sec. 2.2 can effectively be implemented to improve the accuracy of the resulting models. With this aim, we consider a dataset made of pristine hydrogen configurations for which we computed VMC energy, forces, and pressure using different types of WF and with/without bias correction. The dataset is the same as the one used in Ref. Tenti2024 and comprises $558$ configurations of $128$ atoms each. These configurations were extracted from DFT-driven MD simulations at temperatures between $4000$ K and $20000$ K and Wigner-Seitz radii $r_{s}$ between $1.80$ and $2.12$. This range of thermodynamic conditions corresponds to the expected location of the Hugoniot curve of deuterium, defined as the set of states that can be reached by producing a shock wave in a sample. For these configurations, we fully optimized the wave function, going beyond the frozen-orbital JSD ansatz. As previously mentioned, we take this dataset as the reference one. In Sec. 3.1, we study the SCE for the same set of configurations as the reference dataset, and show the effect of both forces and pressure corrections. In Sec. 3.2 we directly compare the accuracy of MLIPs trained on a biased dataset (i.e. affected by the SCE) and a corrected one, respectively. We also compare both models with the reference dataset used in Ref. Tenti2024. Finally, in Sec. 3.3, we analyze the results of molecular dynamics simulations driven by the different MLIPs.